How much gas does it take to mow a whole lawn if it takes 3/4 gallon to mow 6/7 of the whole lawn?
Found 2 solutions by flame8855, GFeliz:
Answer by flame8855(424)

= 3/4 * 7/6 = 7/8 g

Answer by GFeliz(4)

(6/7)/1 = (3/4)/g, where g represents the amount of gas needed
(6/7)g = (3/4)
g = (3/4) divided by (6/7)
g = (3/4) times (7/6)
g = (7/8) gas is needed